OpenStudy (cedric):

simplify 18y^3 over 36xy^2

OpenStudy (anonymous):

cancel common factors from top and bottom... can you do that for me? \[\huge \frac{18y^3}{36xy^2} \]

OpenStudy (cedric):

ok ok uhm like you mean 8 ?

OpenStudy (anonymous):

\[\huge \frac{18y^3}{36xy^2}=\frac{2\cdot 3\cdot 3\cdot y\cdot y\cdot y}{2\cdot 2\cdot 3\cdot 3\cdot x\cdot y\cdot y} \]

OpenStudy (cedric):

oh ok so you'd be left with y over 2x

OpenStudy (anonymous):

yep...:)

OpenStudy (cedric):

medal for you good sir

OpenStudy (anonymous):

thanks...:)

OpenStudy (anonymous):

not sir... though.... :)

OpenStudy (cedric):

oh lol ma'am